# Mutual Information Optimization for SIM-Based Holographic MIMO Systems

### v1

@@ V Computational Complexity
In this section, the computational complexity of 1 is analyzed by counting the required number of complex multiplications. We assume that $N\gg N_{t}$ and $E\gg N_{r}$, which is the usual case for a SIM-based HMIMO communication system. We also assume that all matrices $\APLup\mathbf{x}_{i,j}\APLup\mathbf{x}_{i,j}^{H}$ are precomputed. In the sequel, we provide the complexity for the computation of the precoding matrix $\mathbf{P}$, and the transmit and the receive SIM phase shifts $\{\boldsymbol{\phi}^{l}\}_{l=1}^{L}$ and $\{\boldsymbol{\psi}^{k}\}_{k=1}^{K}$, respectively.
The complexity of calculating all products $e^{-\frac{F_{i,j}}{4\sigma^{2}}}\APLup\mathbf{x}_{i,j}\APLup\mathbf{x}_{i,j}^{H}$ in (13) is $\mathcal{O}(N_{vec}^{2}N_{s}^{2})$, this is also the complexity of calculating the gradient $\nabla_{\mathbf{P}}f(\mathbf{P},\boldsymbol{\phi},\boldsymbol{\psi})$. In addition, $\mathcal{O}(N_{vec}^{2}N_{s}^{2})$ multiplications are required to obtain $f(\mathbf{P}_{n+1},\boldsymbol{\phi}_{n},\boldsymbol{\psi}_{n})$. Hence, the complexity of computing the precoding matrix $\mathbf{P}$ is equal to $\mathcal{O}(I_{p}N_{vec}^{2}N_{s}^{2})$, where $I_{p}$ is the number of line search steps.
The complexity of $\sum_{i,j}e^{-\frac{F_{i,j}}{4\sigma^{2}}}\APLup\mathbf{x}_{i,j}\APLup\mathbf{x}_{i,j}^{H}$ in (14) is $\mathcal{O}(N_{vec}^{2}N_{s}^{2})$. The complexity of calculating the matrix product $\mathbf{G}^{H}\mathbf{Z}^{H}\mathbf{H}\mathbf{P}\sum_{i,j}e^{-\frac{F_{i,j}}{4\sigma^{2}}}\APLup\mathbf{x}_{i,j}\APLup\mathbf{x}_{i,j}^{H}\mathbf{P}^{H}$ can be neglected. Furthermore, $\mathcal{O}(LN^{3})$ multiplications are needed to calculate $\mathbf{L}_{l}$. Hence, the complexity of calculating the gradient $\nabla_{\boldsymbol{\phi}^{l}}f(\mathbf{P},\boldsymbol{\phi},\boldsymbol{\psi})$ is $\mathcal{O}(N_{vec}^{2}N_{s}^{2}+LN^{3})$. After obtaining $\boldsymbol{\phi}_{n+1}^{l}$, the calculation of $\mathbf{B}$ has the complexity of $\mathcal{O}(L\text{$N^{3}$})$ and $\mathcal{O}(MN\min(N_{t},N_{r}))$ multiplications are required to compute $\mathbf{H}$. The complexity of calculating $f(\mathbf{P}_{n+1},\bar{\boldsymbol{\phi}}_{n+1}^{l},\boldsymbol{\psi}_{n})$ is $\mathcal{O}(N_{vec}^{2}N_{s}^{2})$. Therefore, the complexity of computing the transmit SIM phase shifts $\{\boldsymbol{\phi}^{l}\}_{l=1}^{L}$ is given by $\mathcal{O}(L[N_{vec}^{2}N_{s}^{2}+LN^{3}+I_{\boldsymbol{\phi}}(L\text{$N^{3}$}+MN\min(N_{t},N_{r})+N_{vec}^{2}N_{s}^{2})])$, where $I_{\boldsymbol{\phi}}$ is the number of line search loops. In a similar way, the complexity of computing the receive SIM phase shifts $\{\boldsymbol{\psi}^{k}\}_{k=1}^{K}$ is equal to $\mathcal{O}(K[N_{vec}^{2}N_{s}^{2}+KE^{3}+I_{\boldsymbol{\psi}}(K\text{$E^{3}$}+EN\min(N_{t},N_{r})+N_{vec}^{2}N_{s}^{2})])$, where $I_{\boldsymbol{\psi}}$ is the number of line search steps.
@@
$C_{\text{APGM}}$ $=\mathcal{O}(I_{p}N_{vec}^{2}N_{s}^{2}+L[N_{vec}^{2}N_{s}^{2}+LN^{3}+I_{\boldsymbol{\phi}}(L\text{$N^{3}$}+$
$MN\min(N_{t},N_{r})+N_{vec}^{2}N_{s}^{2})]+K[N_{vec}^{2}N_{s}^{2}+KE^{3}$
$+I_{\boldsymbol{\psi}}(K\text{$E^{3}$}+EN\min(N_{t},N_{r})+N_{vec}^{2}N_{s}^{2})]).$ (19)
@@ VI Simulation Results
In this section, we evaluate the CR and the MI of 1 in a SIM-based HMIMO setup. The channel matrix between the transmit and the receive SIM is modeled based on the spatially-correlated channel model as $\mathbf{G}=\mathbf{R}_{\text{R}}^{1/2}\bar{\mathbf{G}}\mathbf{R}_{\text{T}}^{1/2}\in\mathbb{C}^{E\times L}$ [10, 11] where $\bar{\mathbf{G}}\in\mathbb{C}^{E\times L}$ is distributed according to $\mathcal{CN}(0,\beta\mathbf{I})$; $\beta$ is the free space path loss between the transmit and the receive SIM modeled as $\beta(d)=\beta(d_{0})+10b\log_{10}(d/d_{0})$, where $\beta(d_{0})=20\log_{10}(4\pi d_{0}/\lambda)$ is the free space path loss at the reference distance $d_{0}$, $b$ is the path loss exponent, and $d$ is the distance between the transmitter and the receiver. Moreover, $\mathbf{R}_{\text{T}}\in\mathbb{C}^{L\times L}$ and $\mathbf{R}_{\text{R}}\in\mathbb{C}^{E\times E}$ are the spatial correlation matrices of the transmit and the receive SIM, respectively, and the elements of these matrices are given by [10, Eq. (14), (15)].
@@
In the following simulation setup, the parameters are $\lambda=5\,\text{cm}$ (i.e., $f=6\,\text{GHz}$), $N_{t}=2$, $N_{r}=2$, $N_{s}=2$, $\beta=3.5$, $d_{0}=1\,\text{m}$, $d=300\,\text{m}$, $L=K=4$ and $\sigma^{2}=-110\,\text{dB}$. Both the transmit and the receive antennas are placed in arrays parallel to the $x$-axis, and the midpoints of these arrays have coordinates $(0,0,0)$ and $(0,0,d)$, respectively. Also, the inter-antenna separations of these arrays are $\lambda/2$. The meta-atoms in every SIM layer are uniformly placed in a square formation and the size of each meta-atom is $\frac{\lambda}{2}\times\frac{\lambda}{2}$ (i.e., $A=\lambda^{2}/4$). Moreover, all SIM layers are parallel to the $xy$-plane and their centers are along the $z$-axis. The separation between the neighboring SIM layers, as well as the separation between the first SIM layer and the adjacent antenna array are $\lambda/2$. In the line search procedures in 1, all step sizes are initially set to 1000, $\delta=10^{-3}$ and $\rho=1/2$. The initial values of the optimization variables $\mathbf{P}$, $\boldsymbol{\phi}$, and $\boldsymbol{\psi}$ are randomly generated. All results are averaged over 30 independent channel realizations.
@@
In Fig. 1, we present the CR and the MI of the proposed APGM algorithm for different sizes of the discrete symbol alphabet and different numbers of meta-atoms in SIM layers. In general, the MI reflects the same trend as observed in the CR, albeit always achieving larger values. This justifies the use of the CR as an alternative metric for the MI optimization. In all cases, the CR and the MI reach 90 % of their convergent values in approximately 20 iterations of the proposed algorithm. Furthermore, we notice that the change of $M$ has a negligible influence on the CR, and that the MI may even decrease slightly if $M$ increases when the transmit and receive SIMs have 49 meta-atoms per layer. This phenomenon is due to the reduced separation between adjacent constellation points when $M$ is increased without a corresponding increase in the average symbol energy, which consequently increases the bit error probability (BEP) per transmission interval. On the other hand, the CR and the MI demonstrate significant improvements as the number of meta-atoms in the SIM layers increases. This enhancement reaffirms the fact that the beamforming capabilities of SIMs are highly dependent on the number of meta-atoms in SIM layers.
@@
In Fig. 2, we show the MI of the considered system with and without signal precoding. It can be clearly observed that signal precoding can substantially increase the MI for about 47 % and 32 % for 49 and 100 meta-atoms per SIM layer, respectively. On the other side, signal precoding alone, i.e., in the absence of the transmit and the receive SIMs, is only able to provide a near zero MI, which is not shown in Fig. 2. Hence, we can conclude that while digital signal precoding of small scale alone shows limited beamforming gain, but its integration with SIM-based HMIMO systems can generate significant impact on the achievable MI.
@@ VII Conclusion
In this paper, we have demonstrated that the MI in a SIM-based HMIMO system can be efficiently optimized with the channel CR. To maximize the CR, we proposed the APGM which optimizes for the CR by adjusting the transmit precoding, and the phase shifts for the transmit and the receive SIMs in a layer-by-layer basis. Simulation results show that the CR is indeed a reliable metric for optimizing the MI in SIM-based HMIMO systems. Also, integrating even a small scale digital precoder in the considered system can substantially increase the MI performance.
@@
The gradient of $f(\mathbf{P},\boldsymbol{\phi},\boldsymbol{\psi})$ with respect to $\mathbf{P}^{*}$ is given by
$\nabla_{\mathbf{P}}f(\mathbf{P},\boldsymbol{\phi},\boldsymbol{\psi})=-\frac{1}{4\sigma^{2}}\sum_{i,j=1}^{N_{s}}\exp\left(-\frac{F_{i,j}}{4\sigma^{2}}\right)\nabla_{\mathbf{P}}F_{i,j}.$ (20)
@@
The gradient of $f(\mathbf{P},\boldsymbol{\phi},\boldsymbol{\psi})$ with respect to $\boldsymbol{\phi}^{l*}$ is given by
$\nabla_{\boldsymbol{\phi}^{l}}f(\mathbf{P},\boldsymbol{\phi},\boldsymbol{\psi})=-\frac{1}{4\sigma^{2}}\sum_{i,j=1}^{N_{s}}\exp\left(-\frac{F_{i,j}}{4\sigma^{2}}\right)\nabla_{\boldsymbol{\phi}^{l}}F_{i,j}.$ (22)
@@
$\text{d}F_{i,j}$ $=\operatorname{Tr}\{\mathbf{H}\mathbf{P}\APLup\mathbf{x}_{i,j}\APLup\mathbf{x}_{i,j}^{H}\mathbf{P}^{H}\text{d}\mathbf{H}^{H}+\mathbf{P}\APLup\mathbf{x}_{i,j}\APLup\mathbf{x}_{i,j}^{H}\mathbf{P}^{H}\mathbf{H}^{H}\text{d}\mathbf{H}\}$
$=\operatorname{Tr}\{\mathbf{L}_{l}\text{d}(\boldsymbol{\mathbf{\Phi}}^{l})^{H}+\mathbf{L}_{l}^{H}\text{d}\boldsymbol{\mathbf{\Phi}}^{l}\}=\operatorname{Tr}\{\mathbf{L}_{l}^{T}\text{d}\boldsymbol{\mathbf{\Phi}}^{l*}+\mathbf{L}_{l}^{H}\text{d}\boldsymbol{\mathbf{\Phi}}^{l}\}$

### v2

@@ V Computational Complexity
In this section, the computational complexity of 1 is analyzed by counting the required number of complex multiplications. We assume that $N\gg N_{t}$ and $E\gg N_{r}$, which is the usual case for a SIM-based HMIMO communication system. We also assume that all matrices $\APLup\mathbf{x}_{i,j}\APLup\mathbf{x}_{i,j}^{H}$ are precomputed.
The complexity of the gradient $\nabla_{\mathbf{P}}f(\mathbf{P},\boldsymbol{\phi},\boldsymbol{\psi})$ is $\mathcal{O}(N_{vec}^{2}N_{s}^{2})$. In addition, $\mathcal{O}(N_{vec}^{2}N_{r}N_{s})$ multiplications are required to obtain $f(\mathbf{P}_{n+1},\boldsymbol{\phi}_{n},\boldsymbol{\psi}_{n})$. Hence, the complexity of computing the precoding matrix $\mathbf{P}$ is equal to $\mathcal{O}(I_{p}N_{vec}^{2}N_{r}N_{s})$, where $I_{p}$ is the number of line search steps.
The complexity of $\sum_{i,j}e^{-\frac{F_{i,j}}{4\sigma^{2}}}\APLup\mathbf{x}_{i,j}\APLup\mathbf{x}_{i,j}^{H}$ in (11) is $\mathcal{O}(N_{vec}^{2}N_{s}^{2})$. Furthermore, $\mathcal{O}(LN^{3})$ multiplications are needed to calculate $\mathbf{L}_{l}$. Hence, the complexity of calculating the gradient $\nabla_{\boldsymbol{\phi}^{l}}f(\mathbf{P},\boldsymbol{\phi},\boldsymbol{\psi})$ is $\mathcal{O}(N_{vec}^{2}N_{s}^{2}+LN^{3})$. After obtaining $\boldsymbol{\phi}_{n+1}^{l}$, the calculation of $\mathbf{B}$ and $\mathbf{H}$ has the complexity of $\mathcal{O}(\text{$N^{3}$}+ENN_{s})$. The complexity of calculating $f(\mathbf{P}_{n+1},\bar{\boldsymbol{\phi}}_{n+1}^{l},\boldsymbol{\psi}_{n})$ is $\mathcal{O}(N_{vec}^{2}N_{r}N_{s})$. Therefore, the complexity of computing the transmit SIM phase shifts $\{\boldsymbol{\phi}^{l}\}_{l=1}^{L}$ can be approximated as $\mathcal{O}(L[LN^{3}+I_{\boldsymbol{\phi}}(\text{$N^{3}$}+ENN_{s}+N_{vec}^{2}N_{r}N_{s})])$, where $I_{\boldsymbol{\phi}}$ is the number of line search loops. Similarly, the complexity of computing the receive SIM phase shifts $\{\boldsymbol{\psi}^{k}\}_{k=1}^{K}$ is equal to $\mathcal{O}(K[KE^{3}+I_{\boldsymbol{\psi}}(\text{$E^{3}$}+ENN_{s}+N_{vec}^{2}N_{r}N_{s})])$, where $I_{\boldsymbol{\psi}}$ is the number of line search steps.
@@
$C_{\text{APGM}}=\mathcal{O}(I_{p}N_{vec}^{2}N_{r}N_{s}+L[LN^{3}+I_{\boldsymbol{\phi}}(\text{$N^{3}$}+ENN_{s}$
$+N_{vec}^{2}N_{r}N_{s})]+K[KE^{3}+I_{\boldsymbol{\psi}}(\text{$E^{3}$}+ENN_{s}+N_{vec}^{2}N_{r}N_{s})]).$
@@ VI Simulation Results
In this section, we evaluate the CR and the MI of 1 in a SIM-based HMIMO setup. The channel matrix between the transmit and the receive SIM is modeled based on the spatially-correlated channel model as $\mathbf{G}=\mathbf{R}_{\text{R}}^{1/2}\bar{\mathbf{G}}\mathbf{R}_{\text{T}}^{1/2}\in\mathbb{C}^{E\times L}$ [13, 14] where $\bar{\mathbf{G}}\in\mathbb{C}^{E\times L}$ is distributed according to $\mathcal{CN}(0,\beta\mathbf{I})$; $\beta$ is the free space path loss between the transmit and the receive SIM modeled as $\beta(d)=\beta(d_{0})+10b\log_{10}(d/d_{0})$, where $\beta(d_{0})=20\log_{10}(4\pi d_{0}/\lambda)$ is the free space path loss at the reference distance $d_{0}$, $b$ is the path loss exponent, and $d$ is the distance between the transmitter and the receiver. Moreover, $\mathbf{R}_{\text{T}}\in\mathbb{C}^{L\times L}$ and $\mathbf{R}_{\text{R}}\in\mathbb{C}^{E\times E}$ are the spatial correlation matrices of the transmit and the receive SIM, respectively, and their elements are given by [13].
@@
In the following simulation setup, the parameters are $\lambda=5\,\text{cm}$ (i.e., $f=6\,\text{GHz}$), $N_{t}=2$, $N_{r}=2$, $N_{s}=2$, $\beta=3.5$, $d_{0}=1\,\text{m}$, $d=300\,\text{m}$, $L=K=4$ and $\sigma^{2}=-110\,\text{dB}$. Both the transmit and the receive antennas are placed in arrays parallel to the $x$-axis, and the midpoints of these arrays have coordinates $(0,0,0)$ and $(0,0,d)$, respectively. Also, the inter-antenna separations of these arrays are $\lambda/2$. The meta-atoms in every SIM layer are uniformly placed in a square formation and the size of each meta-atom is $\frac{\lambda}{2}\times\frac{\lambda}{2}$. Moreover, all SIM layers are parallel to the $xy$-plane and their centers are along the $z$-axis. The separation between the neighboring SIM layers, as well as the separation between the first SIM layer and the adjacent antenna array are $\lambda/2$. In the line search procedures in 1, all step sizes are initially set to 1000, $\delta=10^{-3}$ and $\rho=1/2$. The initial values of the optimization variables $\mathbf{P}$, $\boldsymbol{\phi}$, and $\boldsymbol{\psi}$ are randomly generated. All results are averaged over 30 independent channel realizations.
@@
In Fig. 1, we present the CR and the MI of the proposed APGM algorithm for different sizes of the discrete symbol alphabet and different numbers of meta-atoms in SIM layers. In general, the MI reflects the same trend as observed in the CR, albeit always achieving larger values. This justifies the use of the CR as an alternative metric for the MI optimization. In all cases, the CR and the MI reach 90 % of their convergent values in approximately 20 iterations of the proposed algorithm. Also, we notice that the change of $M$ has a negligible influence on the CR, and that the MI may even decrease slightly if $M$ increases when the transmit and the receive SIMs have 49 meta-atoms per layer. This phenomenon is due to the reduced separation between adjacent constellation points when $M$ is increased without a corresponding increase in the average symbol energy, which consequently increases the bit error probability (BEP) per transmission interval. On the other hand, the CR and the MI demonstrate significant improvements as the number of meta-atoms in the SIM layers increases. This enhancement reaffirms the fact that the beamforming capabilities of SIMs are highly dependent on the number of meta-atoms in SIM layers. To show the efficacy of the proposed method, we also present the MI obtained using the same step size for all layers in the SIM. As can be seen clearly, the corresponding MI is not able to converge even after 200 iterations and is significantly lower than the MI obtained using different step sizes for each SIM layer as done in our method.
@@
In Fig. 2, we show the MI of the considered system with and without signal precoding. It can be clearly observed that signal precoding can substantially increase the MI for about 47 % and 32 % for 49 and 100 meta-atoms per SIM layer, respectively. On the other side, signal precoding alone, i.e., in the absence of the transmit and the receive SIMs, is only able to provide a near zero MI, which is not shown in Fig. 2. Hence, we can conclude that while digital signal precoding of small scale alone shows limited beamforming gain, but its integration with SIM-based HMIMO systems can generate a significant impact on the achievable MI. We also note that the MI is significantly lower when one of the SIMs employs fixed phase shifts, compared to scenarios where both SIMs are tunable. This reduction in the MI is approximately the same regardless of whether the fixed SIM is at the transmit or the receive end.
@@ VII Conclusion
In this paper, we have demonstrated that the MI in a SIM-based HMIMO system can be efficiently optimized with the channel CR. To maximize the CR, we proposed the APGM which optimizes the CR by adjusting the transmit precoding, and the phase shifts for the transmit and the receive SIMs on a layer-by-layer basis. Simulation results showed that the CR is indeed a reliable metric for optimizing the MI in SIM-based HMIMO systems. Also, integrating even a small scale digital precoder in the considered system could substantially increase the MI performance. For future work, an interesting topic is to study the potential of SIMs in the context of emerging integrated sensing and communication systems. Moreover, as a globally optimal solution has not been reported, it remains to be seen how the performance of the proposed method is compared to the optimal performance.
@@
The gradient of $f(\mathbf{P},\boldsymbol{\phi},\boldsymbol{\psi})$ w.r.t. $\mathbf{P}^{*}$ is given by
$\!\!\!\nabla_{\mathbf{P}}f(\mathbf{P},\boldsymbol{\phi},\boldsymbol{\psi})=-\frac{1}{4\sigma^{2}}\sum\nolimits_{i,j=1}^{N_{s}}\exp\left(-\frac{F_{i,j}}{4\sigma^{2}}\right)\nabla_{\mathbf{P}}F_{i,j}.$ (16)
@@
The gradient of $f(\mathbf{P},\boldsymbol{\phi},\boldsymbol{\psi})$ w.r.t. $\boldsymbol{\phi}^{l*}$ is given by
$\!\!\!\!\!\!\nabla_{\boldsymbol{\phi}^{l}}f(\mathbf{P},\boldsymbol{\phi},\boldsymbol{\psi})=-\frac{1}{4\sigma^{2}}\sum\nolimits_{i,j=1}^{N_{s}}\exp\left(-\frac{F_{i,j}}{4\sigma^{2}}\right)\nabla_{\boldsymbol{\phi}^{l}}F_{i,j}.$ (18)
@@
$\text{d}F_{i,j}$ $=\operatorname{Tr}\{\mathbf{L}_{l}\text{d}(\boldsymbol{\mathbf{\Phi}}^{l})^{H}+\mathbf{L}_{l}^{H}\text{d}\boldsymbol{\mathbf{\Phi}}^{l}\}=\operatorname{Tr}\{\mathbf{L}_{l}^{T}\text{d}\boldsymbol{\mathbf{\Phi}}^{l*}+\mathbf{L}_{l}^{H}\text{d}\boldsymbol{\mathbf{\Phi}}^{l}\}$